opcorns:

A sample that weighs 101.75 g is a mixture of 30% helium atoms and 70% krypton atoms.

Vocaloid:

Full question: (please post this next time) A sample that weighs 101.75 gg is a mixture of 30% helium atoms and 70% krypton atoms. Part A How many particles are present in the sample? Express your answer using two significant figures

Razor:

The given sample is 101.75 gg(giga) in a mixture of 30% helium atoms and 70% krypton atoms. First , the mole concept. 1 mole of an element has 6.022 x 10^23 atoms or particles. 1 mole of helium has 4 grams. 4 grams of helium has again - 6.022 x 10^23 atoms or particles. 1 mole of krypton has 84 grams. 84 grams of this element has 6.022 x 10^23 atoms or particles. In order to find the mass of 30% helium atoms and 70% krypton atoms: \[\frac{ 30 }{ 100} \] x 101.75. That comes out as 30.52 grams. \[\frac{ 70 }{ 100}\] x 101.75. That comes out as 71.22 grams. Now that we have the mass... 30.52g of helium element contains: \[\frac{ 6.022 x 10^{23} }{ 4 }\] x 30.52 = 4.594786 or 45.946 x 10^{23} atoms or particles. 71.22 grams of this is \[\frac{6.022 x 10^{23} }{ 84 }\] x 71.22 = 5.1057957 or 51.057 x \[10^{23}\] atoms or particles. (Add the sum from helium and krypton) (45.946 + 51.057) x \[10^{23}\] = 9.6803 or 96.90 x \[10^{23}\] atoms or particles

Razor:

Apologies if the answer comes out a bit janky - I'm a bit rusty with the equation tool. Ahah...

Razor:

Welcome to QuestionCove nonetheless.

Shadow:

@dude